Howard Electronics, a small manufacturer of electronic research equipment, has approximately 7000 items in inventory and has hired Eng. Mostafa to manage its inventory. Mostafa has determined that 10% of the items in inventory are A items, 35% are B items, and 55% are C items. He would like to set up a system in which all A items are counted monthly (every 20 working days), and all B items are counted quarterly (every 60 working days), and all C items are counted semiannually (every 120 working days). How many items need to be counted each day?

Let, each A item needs 10 minutes to be counted, each B item needs 5 minutes, and each B item needs 2 minutes. If the company works one shift 8 hours per day, how many labors needed to count per day?
Answer #1

A items are 0.1 * 7000 = 700

B items are 0.35 * 7000 = 2450

C Items are 0.55 * 700 = 3850

Units to be counted everyday are 700/20 + 2450/60 + 3850/120 = 107.92 or 108 items


answered by: Pushkin